ALGORITHM ENGINEERING HOMEWORK

Let g(n) = n. Is it true that g(n) 2 O(n2)? Why or why not?
Answer #1
f(n) = O(g(n)) means there are positive constants c and n0, such that 0 ≤ f(n) ≤ cg(n) for all n ≥ n0

g(n) = O(n^2)
n    = O(n^2)

=>  n <= c*n^2

Let's assume c = 1

=>  n <= c*n^2
=>  n <= n^2
=>  0 <= n

so, g(n) = O(n^2) given c = 1 and n0 = 0 

Please let me know if you need me to make any changes.